OpenStudy (anonymous):

zWrite the equation of the given circle. center (3, 3) radius of 2

OpenStudy (amoodarya):

circle with center (a,b) radius of r is (x-a)^2+(y-b)^2=r^2

OpenStudy (amoodarya):

so (x-3)^2+(y-3)^2=2^2